Hobart 26m

Disk VSN: HOB+0101/16000/1024 | Data volume at beginning: 10995.1 GB

* note: couldn't flogit? “can't open file” after “Appending comments to log”. Did the upload of log and prc manually (Lim)

  • 1800 UTC Experiment started ok. (Jasper)

* 20:04 UTC Wind Stow caused the telescope so get stuck miss reporting drive status, elevation, azimuth and source. At 20:30 UTC schedule is halted but antenna=open and antenna=operate fail to work. Called Warren who rebooted the field system and went to check the dish onsite. Warren says the dish parked ok originally but to leave the schedule halted until shift change since the wind is still high.

* 2352 - Schedule resumed after leaving windstow, drive PC reboot. First good data from 2352 (JMc) * 0040 - Windstowed again (JMc) * 0220 - Moving again. Had briefly unstowed previously (JMc)